### v1

@@ 4.4 Estimation and inference for the bounds
The above bounds are all functions of the population probabilities $p_{ym\cdot a}$, but researchers only have access to the observed data $\{(A_{i},M_{i},Y_{i},\mathbf{X}_{i})\}_{i=1,\ldots,n}$. To perform estimation and inference for these bounds, we rely on the intersection bounds approach of Chernozhukov, Lee and Rosen (2013). Traditional plug-in estimators for the bounds tend to be biased so that the plug-in bounds are narrower than the population bounds because of the use of the minimum and maximum operators. Furthermore, the estimated standard error of the selected bounding expression would ignore the uncertainty that the other bounding expressions also have, which contributes to uncertainty about the overall bound. The approach of Chernozhukov, Lee and Rosen (2013) uses precision-corrected versions of each bounding expression to obtain (a) upper and lower bound estimates that are half-median unbiased and (b) provides confidence intervals for the true indirect effect that cover the true value of the indirect effect at nominal rates. Half-median unbiased means that the probability that the estimated upper (lower) bound is greater (less) than the population upper (lower) bound is at least one-half. This is a common desirable property of estimators in partially identified settings like ours where previous work has shown that traditional unbiasedness is not possible to achieve (Hirano and Porter, 2012). In the next section, we implement these empirical procedures for our two motivating examples.
@@ 5.1 Reducing Outgroup Prejudice
We first re-analyze data from the Kalla and Broockman (2023) study of intergroup prejudice. In one of their studies, the authors showed respondents a picture of an outgroup member (either undocumented immigrants or transgender people) and then asked them to engage in a randomly assigned exercise corresponding to different narrative strategies. [^5] These are traditional perspective-taking, analogic perspective-taking, perspective getting with an essay task to summarize the story heard, and perspective getting without the essay task. We focus on the pooled (i.e., across the unauthorized immigrants and transgender people arms) effects of the “perspective getting with essay” treatment relative to the baseline condition, since this treatment yields the highest effect on prejudice reduction of the four treatment conditions considered. In this treatment condition, respondents read a story about an outgroup member and were then asked to summarize the story they just read.
@@
To assess which mechanisms explain these effects, the authors asked respondents four additional questions as potential mediators: (i) whether respondents have a lot in common with the outgroup; (ii) whether the suffering of the outgroup is of concern ; (iii) whether outgroup members face challenges that are no fault of their own, and (iv) whether the outgroup’s name elicits thoughts about individual people or a group of people. Importantly for our purposes, the authors implement an IOT approach on these four mediators.
@@
Since our sharp bounds for the indirect effect (as described in Section 4.3) rely on dichotomous mediators and outcomes, we first dichotomize the four mediators and the outcome variable at their medians. We then implement the IOT approach, estimating IOT effects without covariate controls. We do so because the treatment is randomly assigned in this study. As shown in the right panel of Figure 2, these IOT estimates largely replicate the direction, magnitude, and statistical significance of those reported by Kalla and Broockman (2023): perspective-getting with an essay task has a positive and statistically significant effect on whether the suffering of the outgroup is of concern (0.04, 95% confidence interval of [0.03, 0.06]) and whether respondents have a lot in common with the outgroup (also 0.04, 95% confidence interval of [0.03, 0.06]). [^6] Note that the authors also find a positive effect of perspective-getting with an essay task on whether outgroup members face challenges that are no fault of their own, a finding that does not replicate without controlling for the battery of covariates included in the authors’ models. The authors rule out the last of these mediators because placebo interventions also induced a sense of commonality with the outgroup. Thus, they argue that concern about outgroup suffering is likely a key pathway for the effect of their intervention. [^7] Of course, these different mediators overlap to some degree. We treat each mediator separately, but a fruitful path for future research would be to incorporate multiple mediators into these bounds.
@@
Does this conclusion hold if we implement our nonparametric bounds for the indirect effect? Figure 2, right panel, suggests it does not. This panel displays the sharp bounds for the indirect effect with the treatment reference category of $A=0$, i.e., $\delta(0)$, and calibrates these bounds both against zero (dotted black line) and the overall ATE of the intervention (dotted red line), which was $0.03$.
@@
Figure 2: Establishing causal mechanisms for the effect of perspective-getting on prejudice reduction (Kalla and Broockman, 2023). The figure displays IOT point estimates (left panel) and sharp bounds estimates (thick, transparent lines) for the indirect effect of perspective-getting with an essay assignment on prejudice reduction via each of the putative mediators at reference level $A=0$ ($\delta(0)$), as well as the 95% confidence intervals for these bounds (thin lines). No-assumption bounds are shown in blue, while MMR bounds derived under Assumption 3) are shown in green. The dotted black line marks zero, while the dotted red line marks the ATE (0.03).
@@
For both the no-assumption bounds (blue-colored bounds) and the MMR bounds (green-colored bounds), we present the sharp bounds (thick, transparent lines) in addition to their 95% confidence intervals (thin lines). The estimated no-assumption bounds are extremely wide and approximately center around 00, providing little to no information about the existence of an indirect effect for each of the mediators considered, nor their direction or magnitude. In particular, the no-assumption bounds estimated for the key mediator cited by Kalla and Broockman (2023) – reactive empathy – is $[-0.39,0.50]$, an interval that contains a large range of positive and negative effect sizes, contains zero, and includes a set of indirect effect sizes on the order of 10 times larger in absolute value than the estimate recovered by the IOT approach. In short, the no-assumption bounds do not allow us to conclude that any of the four mediators considered mediate the effect of perspective-getting on prejudice reduction, even those mediators that “pass” an IOT (i.e., those mediators that are themselves affected by the treatment), such as reactive empathy and attributional thinking.
@@
Does assuming mediator monotonicity – that the treatment affects the mediator in one direction for all respondents – enable us to recover the informativeness of the IOT approach? The green bands in Figure 2 panel B address this question. Two points are of particular note. First, the bounds shrink substantially compared with those computed without restriction on the distribution of potential outcomes. For example, the bounds estimate for reactive empathy shrinks from $[-0.39,0.50]$ under no assumption to $[-0.05,0.043]$ under MMR, a considerably smaller interval that ranges from approximately $-ATM$ to $+ATM$. Second, and despite this, for the two putative mediators with non-zero IOT estimates (concern about outgroup suffering and feeling commonality with the outgroup), the bounds are uninformative about either the existence or the sign of the relevant indirect effect. The bounds all contain zero and a large range of positive and negative values, including the overall ATE (red dotted line). In other words, the observed data are consistent with the indirect effect via these mediators being zero, as well as with the indirect effect via these mediators being equal to the ATE (and thus, consistent with these mediators serving as the sole mechanism through which the ATE operates). Only for the group vs. individual mediator – and to a lesser extent, the “outgroup suffering is not their fault” mediator – is the MMR bounds estimate extremely narrow: in this case, it almost collapses on zero, a result which mirrors the IOT finding of an extremely small effect of treatment on those mediators. This result highlights the ‘approximately necessary” condition of IOTs under MMR, in that small ATMs imply small indirect effects.
Is the MMR a reasonable assumption in this setting? For the group vs. individual mediator proposed by Kalla and Broockman (2023), this means assuming that no individual for whom hearing about the feelings of prejudice of trans people or immigrants would make them more likely to view them as individuals rather than as a group. Given the amorphous nature of this kind of evaluation, it seems difficult to maintain this strong of an assumption.
@@ 5.2 Transitional Justice Museums and Support for Democracy
The authors posit that museums’ narratives of past events may have elicited emotional reactions among their visitors. To test this hypothesis, they measured treated and untreated students’ responses to the Positive and Negative Affect Schedule, which asks respondents to consider a range of positive and negative emotions and to indicate which they feel in the present moment (Watson, Clark and Tellegen, 1988). We focus on the negative emotion questions as the authors find treatment effects on these emotions to be substantively larger and more consistent than the positive emotion questions. Similar to the previous example, we dichotomize all mediators as well as the outcome variable at their medians. Moreover, given the randomized nature of treatment in this study, we estimate the IOT effects using bivariate regression between treatment and mediator, similar to the previous application. The left panel of Figure 3 reports the IOT estimates, which replicate the patterns reported by Balcells, Palanza and Voytas (2022): the visit to the transitional justice museum has a positive and statistically significant effect on the overall negative emotions experienced by respondents (0.46, 95% confidence interval of $[0.35,0.57]$). In particular, the authors find positive and significant treatment effects on feeling tense, scared, guilty, hostile, fearful, afraid, and embarrassed. Balcells, Palanza and Voytas (2022) argue that these findings indicate that the emotional experience of a transitional justice museum may act as a key mechanism by which the museum visit alters political attitudes.
@@
Figure 3: Establishing causal mechanisms for the effect of a transitional justice museum visit on support for democratic institutions (Balcells, Palanza and Voytas, 2022). The figure displays IOT point estimates (left panel) and sharp bounds estimates (thick, transparent lines) for the indirect effect of the museum intervention via each of the putative mediators at reference level $A=0$ ($\delta(0)$), as well as the 95% confidence intervals for these bounds (thin lines). No-assumption bounds are shown in blue, while MMR bounds derived under Assumption 3) are shown in green. The dotted black line indicates zero, the dotted red line indicates the ATE (0.15) of the museum visit on support for democracy.
@@
While the IOTs could indicate that the strong and significant emotional reactions elicited by a museum visit could serve as an important mediating pathway for the overall museum effect, the right panel of fig. 3 cautions against such an interpretation. Paralleling the canvassing example, we make three remarks about the bounds on $\delta(0)$. First, all no-assumption bounds are extremely wide, and all contain zero. For example, the bounds for the overall negative emotion outcome is [-0.22, 0.80]. Second, the MMR bounds are significantly shorter than their no-assumption counterparts, although they still contain zero. Therefore, while the MMR assumption seems relatively plausible under this empirical setting—that compared to the control condition, visiting a museum recounting the country’s past political violence only increases negative emotions such as fear and guilt—the bounds derived under this assumption remain uninformative about mediation via each of the variables considered. Finally, smaller IOT coefficients correspond to shorter bounds under the MMR assumption. For instance, the effect of the museum visit on “feeling irritable” has a point estimate of 0.07, the smallest among all estimates. Correspondingly, the bound for “feeling irritable” is also the shortest, ranging from $[-0.09,0.07]$, which is approximately $-ATM$ to $+ATM$. Even for these mediators, however, the confidence intervals are close to the overall ATE, meaning we cannot rule out the possibility that they account for the entire effect of the museum intervention.
@@ 6 What Is the Point of Mechanism Tests?
In the social sciences, the idealized version of empirical research happens by a process where researchers (a) develop a relatively broad causal theory (or explanation), (b) use this theory to derive specific hypotheses about specific causal effects, and (c) use convincing causal identification and suitable data and estimators to test these hypotheses. One hopes the connection between a hypothesis and the estimated causal effect will be strong, but one hypothesis (and the resulting causal effect) may be consistent with several broader causal theories. Researchers hope to use mechanism tests to bridge this gap by showing the path by which the treatment affects the outcome. A causal mechanism helps generalize a specific causal effect into a more general form, allowing us to hopefully better understand all situations in which this mechanism is activated. Unfortunately, there are many instances where all approaches to establishing indirect effects, including IOTs, rely on shaky assumptions that are difficult to justify. What can be done in that case?
We encourage researchers to pursue the time-tested approach of “implication analysis” (Lieberson and Horwich, 2008) or “inference to the best explanation” (Spirling and Stewart, 2024), whereby researchers generate many different hypotheses that differentiate between the proposed theory and alternatives. The benefit of this approach is that these secondary hypotheses can be completely unrelated to the direct or indirect effects of the main causal effect of interest. This can free researchers from focusing on research designs with assumptions ill-suited to a particular setting. Of course, as we have articulated here, we must be clear about the assumptions needed to draw conclusions from these tests. In many ways, the main goal of Kalla and Broockman (2023) takes this exact approach. Earlier work showed that interpersonal conversations can lead to reductions in exclusionary attitudes, but it was not clear why these conversations worked. The experimental design of Kalla and Broockman (2023) is tailored to test different types of narrative strategies to help adjudicate between different theoretical mechanisms of different narrative strategies. Thus, their main analysis helps clarify causal mechanisms without resorting to IOTs, effect heterogeneity, or causal mediation.

### v2

@@ 4.6 Estimation and inference for the bounds
The above bounds are all functions of the population probabilities $p_{ym\cdot a}$, but researchers only have access to the observed data $\{(A_{i},M_{i},Y_{i},\mathbf{X}_{i})\}_{i=1,\ldots,n}$. To perform estimation and inference for these bounds, we rely on the intersection bounds approach of Chernozhukov, Lee and Rosen (2013). Traditional plug-in estimators for the bounds tend to be biased so that the plug-in bounds are narrower than the population bounds because of the use of the minimum and maximum operators. Furthermore, the estimated standard error of the selected bounding expression would ignore the uncertainty that the other bounding expressions also have, which contributes to uncertainty about the overall bound. The approach of Chernozhukov, Lee and Rosen (2013) uses precision-corrected versions of each bounding expression to obtain (a) upper and lower bound estimates that are half-median unbiased and (b) provides confidence intervals for the true indirect effect that cover the true value of the indirect effect at nominal rates. Half-median unbiased means that the probability that the estimated upper (lower) bound is greater (less) than the population upper (lower) bound is at least one-half. This is a common desirable property of estimators in partially identified settings like ours where previous work has shown that traditional unbiasedness is not possible to achieve (Hirano and Porter, 2012). We implement these empirical procedures in the next section for our two motivating examples.
@@ 5.1 Reducing Outgroup Prejudice
We first re-analyze data from the Kalla and Broockman (2023) study of intergroup prejudice. In one of their experimental studies, the authors showed respondents a picture of an outgroup member (either undocumented immigrants or transgender people) and then asked them to engage in a randomly assigned exercise corresponding to different narrative strategies. [^5] These are traditional perspective-taking, analogic perspective-taking, perspective getting with an essay task to summarize the story heard, and perspective getting without the essay task. We focus on the pooled (i.e., across the unauthorized immigrants and transgender people arms) effects of the “perspective getting with essay” treatment relative to the baseline condition since this treatment yields the highest effect on prejudice reduction of the four treatment conditions considered. In this treatment condition, respondents read a story about an outgroup member and were then asked to summarize the story they just read. Overall, the authors find that the treatment significantly reduces respondents’ prejudice against outgroup members.
@@
To assess which mechanisms explain these effects, the authors asked respondents four additional questions as potential mediators: (i) whether respondents have a lot in common with the outgroup (self-outgroup merging); (ii) whether the suffering of the outgroup is of concern (reactive empathy); (iii) whether outgroup members face challenges that are no fault of their own (attributional thinking), and (iv) whether the outgroup’s name elicits thoughts about individual people or a group of people (entativity). The authors implement four separate IOTs on these four mediators.
@@
Since our sharp bounds for the indirect effect (as described in Section 4.5) rely on dichotomous mediators and outcomes, we first dichotomize the four mediators and the outcome variable at their medians. We then implement the IOT by estimating ATM effects for each mediator. [^6] We opt not to use covariate adjustment for these ATM effects, which is valid given the random treatment assignment. As shown in the right panel of Figure 2, these IOT estimates largely replicate the direction, magnitude, and statistical significance of those reported by Kalla and Broockman (2023): perspective-getting with an essay task has a positive and statistically significant effect on whether the suffering of the outgroup is of concern (0.04, 95% confidence interval of [0.03, 0.06]) and whether respondents have a lot in common with the outgroup (also 0.04, 95% confidence interval of [0.03, 0.06]). [^7] Note that the authors also find a positive effect of perspective-getting with an essay task on whether outgroup members face challenges that are no fault of their own, a finding that does not replicate without controlling for the battery of covariates included in the authors’ models after dichotomization.
In their original analysis, Kalla and Broockman (2023) provide a nuanced interpretation of the potential mechanisms. For example, the experiment contains two placebo treatments with null effects on the main outcomes (that is, null estimated ATEs) but significant effects on two potential mediators (attributional thinking and self-outgroup merging). The authors argue that this combination implies that these two mediators are unlikely to explain the ATE of the primary perspective-getting treatment. Combined with the IOT evidence on concern about outgroup suffering, the authors note that “perspective-getting reduces prejudice by prompting empathy towards that group or through other unmeasured mediators” and ultimately conclude that unmeasured mediators may be the more important of these two. Thus, the authors lean more toward the falsification interpretation of IOTs rather than the establishment interpretation but ultimately cast doubt on how much can be learned from the IOT analysis.
@@
What conclusions can we reach with our nonparametric bounding analysis? The right panel of Figure 2 displays the sharp bounds for the indirect effect with the treatment reference category of $A=0$, i.e., $\delta(0)$, and calibrates these bounds both against zero (dotted black line) and the overall ATE of the intervention (dotted red line), which was $0.03$.
@@
Figure 2: Establishing causal mechanisms for the effect of perspective-getting on prejudice reduction (Kalla and Broockman, 2023). The figure displays IOT point estimates (left panel) and sharp bounds estimates (thick, transparent lines) for the indirect effect of perspective-getting with an essay assignment on prejudice reduction via each of the putative mediators at reference level $A=0$ ($\delta(0)$), as well as the 95% confidence intervals for these bounds (thin lines). Randomization bounds are shown in blue, while MMR bounds derived under Assumption 3) are shown in green. The dotted black line marks zero, while the dotted red line marks the ATE (0.03).
@@
For both the randomization bounds (wider blue-colored bounds) and the MMR bounds (narrower green-colored bounds), we present the sharp bounds (thick, transparent lines) in addition to their 95% confidence intervals (thin lines). The estimated randomization bounds are extremely wide and approximately center around 00, providing little to no information about the existence of an indirect effect for each of the mediators considered, nor their direction or magnitude. In particular, the randomization bounds estimated for reactive empathy—a potentially important mediator based on the IOT test—is $[-0.39,0.50]$, an interval that contains a large range of positive and negative effect sizes, contains zero, and includes a set of indirect effect sizes on the order of 10 times larger in absolute value than the estimate recovered by the IOT approach. In short, the randomization bounds do not allow us to conclude that any of the four mediators considered mediate the effect of perspective-getting on prejudice reduction, even those mediators that “pass” an IOT (i.e., those mediators that are themselves affected by the treatment), such as reactive empathy and attributional thinking.
@@
Does assuming mediator monotonicity – that the treatment affects the mediator in one direction for all respondents – enable us to recover the informativeness of the IOT approach? The green bands in Figure 2 panel B address this question. Two points are of particular note. First, the bounds shrink substantially compared with those computed without restriction on the distribution of potential outcomes. For example, the bounds estimate for reactive empathy shrinks from $[-0.39,0.50]$ under no assumption to $[-0.05,0.043]$ under MMR, a considerably smaller interval that ranges from approximately $-ATM$ to $+ATM$.
Second, and despite this, for the two putative mediators with nonzero IOT estimates (concern about outgroup suffering and feeling commonality with the outgroup), the bounds are uninformative about either the existence or the sign of the relevant indirect effect. The bounds all contain zero and a large range of positive and negative values, including the overall ATE (red dotted line). In other words, the observed data are consistent with the indirect effect via these mediators being zero, as well as with the indirect effect via these mediators being equal to the ATE (and thus, consistent with these mediators serving as the sole mechanism through which the ATE operates). Only for the group vs. individual mediator – and to a lesser extent, the “outgroup suffering is not their fault” mediator – is the MMR bounds estimate extremely narrow: in this case, it almost collapses on zero, a result which mirrors the IOT finding of a minimal effect of treatment on those mediators. This result highlights the “approximately necessary” condition of IOTs under MMR, in that small ATMs imply small indirect effects.
Figure 3: Validating the MMR Assumption for the effect of perspective-getting on prejudice reduction (Kalla and Broockman, 2023). The figure displays a histogram for estimated CATE values for the effect of $A_{i}$ on $M_{i}$ for each of the four putative mediators. CATEs are estimated using causal forests (Wager and Athey, 2018) with 2000 trees.
Is the MMR a reasonable assumption in this setting? At a theoretical level, for the group vs. individual mediator proposed by Kalla and Broockman (2023), this means assuming that no individual for whom hearing about the feelings of prejudice of trans people or immigrants would make them more likely to view them as a group rather than as individuals. Given the amorphous nature of this kind of evaluation, we think it is difficult to maintain this strong assumption in the absence of additional evidence. Empirically, we can implement our falsification test of MMR, as discussed in Section  4.5. Specifically, we estimate CATEs ($\mathbb{E}[M_{i}(1)-M_{i}(0)|X_{i}=x]$) using causal forests. This flexible machine-learning method facilitates the discovery of heterogeneous treatment effects that may be highly non-linear across covariate values (Wager and Athey, 2018). This nonparametric approach is particularly advantageous in our context, where our falsification goal is to see whether any CATE falls below zero—capturing CATEs flexibly without imposing any particular parametric functional form allows us to assess this claim honestly. Figure 3 plots histograms for each of the four putative mediators. We see that in all cases, a non-trivial proportion of estimated CATEs falls below zero. In this way, MMR appears to fail for all mediators in this empirical setting. Thus, in this particular case, we believe the uninformative randomization bounds are more appropriate, which is in line with the authors’ inconclusive interpretation of mechanisms in the original paper.
@@ 5.2 Transitional Justice Museums and Support for Democracy
The authors posit that museums’ narratives of past events may have elicited emotional reactions among their visitors. To test this hypothesis, they measured treated and untreated students’ responses to the Positive and Negative Affect Schedule, which asks respondents to consider a range of positive and negative emotions and to indicate which they feel in the present moment (Watson, Clark and Tellegen, 1988). We focus on the negative emotion questions as the authors find treatment effects on these emotions to be substantively larger and more consistent than the positive emotion questions. Similar to the previous example, we dichotomize all mediators as well as the outcome variable at their medians. The left panel of Figure 4 reports the IOT estimates, which replicate the patterns reported by Balcells, Palanza and Voytas (2022): the visit to the transitional justice museum has a positive and statistically significant effect on the overall negative emotions experienced by respondents (0.46, 95% confidence interval of $[0.35,0.57]$). In particular, the authors find positive and significant treatment effects on feeling tense, scared, guilty, hostile, fearful, afraid, and embarrassed. Drawing on the establishment interpretation of IOTs, Balcells, Palanza and Voytas (2022) argue that these findings indicate that the emotional experience of a transitional justice museum may act as a key mechanism by which the museum visit alters political attitudes.
@@
Figure 4: Establishing causal mechanisms for the effect of a transitional justice museum visit on support for democratic institutions (Balcells, Palanza and Voytas, 2022). The figure displays IOT point estimates (left panel) and sharp bounds estimates (thick, transparent lines) for the indirect effect of the museum intervention via each of the putative mediators at reference level $A=0$ ($\delta(0)$), as well as the 95% confidence intervals for these bounds (thin lines). Randomization bounds are shown in blue, while MMR bounds derived under Assumption 3) are shown in green. The dotted black line indicates zero; the dotted red line indicates the ATE (0.15) of the museum visit on support for democracy.
@@
While the IOTs could indicate that the strong and significant emotional reactions elicited by a museum visit could serve as an important mediating pathway for the overall museum effect, the nonparametric bounds in the right panel of Figure 4 provide a more inconclusive interpretation. Paralleling the previous canvassing example, we make three remarks about the bounds on $\delta(0)$. First, all randomization bounds are extremely wide, and all contain zero. For example, the bounds for the overall negative emotion outcome is $[-0.22,0.80]$, which confirms our proposition that randomization alone can neither establish nor rule out the existence of mechanisms. Second, the MMR bounds are significantly shorter than their randomization counterparts, although they still contain zero. Therefore, while the MMR assumption seems relatively plausible under this experimental setting—that compared to the control condition, visiting a museum recounting the country’s past political violence only increases negative emotions such as fear and guilt—the bounds derived under this assumption remain uninformative about mediation via each of the variables considered. Moreover, as we show in the Supplemental Materials (Figure SM.5), the estimated CATEs from a causal forest analysis for each of the individual emotions do not fall below zero. Finally, smaller IOT coefficients correspond to shorter bounds under the MMR assumption. For instance, the effect of the museum visit on “feeling irritable” has a point estimate of 0.07, the smallest among all estimates. Correspondingly, the bound for “feeling irritable” is also the shortest, ranging from $[-0.09,0.07]$, which is approximately $-ATM$ to $+ATM$. Even for these mediators, however, the confidence intervals are close to the overall ATE, meaning we cannot rule out the possibility that they account for the entire effect of the museum intervention.
@@ 6 What Is the Point of Mechanism Tests?
In the social sciences, the idealized version of empirical research happens by a process where researchers (a) develop a relatively broad causal theory (or explanation), (b) use this theory to derive specific hypotheses about specific causal effects, and (c) use convincing causal identification and suitable data and estimators to test these hypotheses. One hopes the connection between a hypothesis and the estimated causal effect will be strong, but one hypothesis (and the resulting causal effect) may be consistent with several broader causal theories. Researchers hope to use mechanism tests to bridge this gap by showing the path by which the treatment affects the outcome. A causal mechanism helps generalize a specific causal effect into a more general form, allowing us to hopefully better understand all situations in which this mechanism is activated. Unfortunately, there are many instances where all approaches to establishing indirect effects, including IOTs, rely on shaky assumptions that are difficult to justify. What should researchers do in that case?
We encourage researchers to pursue the time-tested approach of “implication analysis” (Lieberson and Horwich, 2008) or “inference to the best explanation” (Spirling and Stewart, 2024), whereby researchers generate many different hypotheses that differentiate between the proposed theory and alternatives. The benefit of this approach is that these secondary hypotheses can be completely unrelated to the direct or indirect effects of the main causal effect of interest. This approach can free researchers from focusing on research designs with assumptions ill-suited to a particular setting. Of course, as we have articulated here, we must be clear about the assumptions needed to draw conclusions from these tests. In many ways, the main goal of Kalla and Broockman (2023) takes this exact approach. Earlier work showed that interpersonal conversations can lead to reductions in exclusionary attitudes, but it was not clear why these conversations worked. The experimental design of Kalla and Broockman (2023) is tailored to test different types of narrative strategies to help adjudicate between different theoretical mechanisms of different narrative strategies. Thus, their primary analysis helps clarify causal mechanisms without resorting to IOTs, effect heterogeneity, or causal mediation.